# Enhancing Textual Personality Detection toward Social Media: Integrating Long-term and Short-term Perspectives

### v1

@@ Abstract
Textual personality detection aims to identify personality characteristics by analyzing user-generated content toward social media platforms. Numerous psychological literature highlighted that personality encompasses both long-term stable traits and short-term dynamic states. However, existing studies often concentrate only on either long-term or short-term personality representations, without effectively combining both aspects. This limitation hinders a comprehensive understanding of individuals’ personalities, as both stable traits and dynamic states are vital. To bridge this gap, we propose a Dual Enhanced Network (DEN) to jointly model users’ long-term and short-term personality for textual personality detection. In DEN, a Long-term Personality Encoding is devised to effectively model long-term stable personality traits. Short-term Personality Encoding is presented to capture short-term dynamic personality states. The Bi-directional Interaction component facilitates the integration of both personality aspects, allowing for a comprehensive representation of the user’s personality. Experimental results on two personality detection datasets demonstrate the effectiveness of the DEN model and the benefits of considering both the dynamic and stable nature of personality characteristics for textual personality detection.
keywords:
personality detection , personality development , long-term stable traits , short-term dynamic states
[…]
1 Introduction
Textual personality detection on scocial media aims to identify an individual’s personality characteristics as implied in the text they created [32]. By analyzing a user’s personality characteristics, we can gain a deeper understanding of their preferences [9], behavior patterns [25], and decision-making processes [10]. As a result, personality detection has been widely applied in various downstream tasks, such as personalized advertising and recommendation systems [6], dialogue systems [31], and individual career planning [14].
@@ 1 Introduction
Traditionally, personality assessment relies on questionnaires designed by psychologists, which can be time-consuming and labor-intensive, and impractical for the online world. The rise of social media has created vast opportunities for automated textual personality detection, attracting a growing number of researchers in the field [8]. Textual personality detection, although resembling a multi-label classification task, poses unique challenges. For example, the input comprises multiple posts rather than a single text [32]. Each post can reflect the user’s current short-term personality state, while the collection of all posts can capture the user’s long-term trait. To address these challenges, numerous studies have been conducted, broadly categorized into two main types.
On the one hand, some methods utilize a flat structure and focus on modeling the overall long-term representation of users to capture their stable personality traits for personality detection. This can be accomplished by utilizing various linguistic or statistical features [23, 35] or concatenating the user’s post representations into a single representation [11] to capture the long-term representation. On the other hand, another line of research, employing a hierarchical structure, involves first modeling the current short-term representation of users’ states using their individual posts. Subsequently, these representations are aggregated to obtain an overall user representation for personality detection [18, 30, 32].
@@
Figure 2: Distribution of changes in MBTI personality characteristics among 84 college students throughout their college years. The bar chart in the figure compares the number of personality changes among different gender groups, while the line graph illustrates the trend of the percentage of different personality change counts.
@@
The above findings emphasize the complexity of personality development, with a combination of stability and change occurring simultaneously. It underscores the importance of considering both the stable long-term traits and the dynamic short-term states when studying and predicting personality. Unfortunately, it is non-trivial to capture both the long-term and short-term perspectives of user personality due to the following obstacles: (1) How to model the stable, long-term personality traits of users? (2) How to model the dynamic changes in users’ short-term personality states? (3) How to dynamically interact between these two personality representations to obtain a relatively stable user personality for personality detection?
To address the three key challenges, we propose the Dual Enhanced Network (DEN) for personality detection that captures both long-term and short-term personality representations, dynamically integrating them for effective personality detection. Specificly, DEN, introduces three novel modules that jointly model users’ long-term and short-term personality characteristics and dynamically interact between them: (1) Long-term Personality Encoding module utilizes a graph neural network (GNN) based on the psychological graph constructed using LIWC(Linguistic Inquiry and Word Count) [23] to extract users’ long-term personality traits. By employing the GNN on the user psychological graph, it effectively analyzes the usage patterns of psychological entities and learns the long-term personality representations of users, providing valuable insights into their enduring characteristics; (2) Short-term Personality Encoding module utilizes a pretrained language model that employs the multi-head self-attention mechanism and excels in contextualizing text and to dynamically capture users’ short-term personality changes, enabling effective modeling of short-term states; and (3) Bi-directional Interaction module dynamically constructs a novel bipartite graph that facilitates interaction between long-term and short-term representations, enhancing both representations. By integrating these enhanced representations, the model can make more accurate predictions of personality characteristics. Extensive experiments demonstrate the effectiveness of DEN in the personality detection task. The key contributions of this paper can be summarized as follows:
@@ 2 Related Work
One type of research in textual personality detection utilizes a flat structure, focusing on modeling the overall long-term representation of users, which reflects their stable personality traits. Various approaches have been proposed in this category. For instance, studies such as [23, 19, 26, 21] employ various linguistic features extracted from the entirety of users’ posts to capture their language habits and represent their long-term traits. Similarly, [35, 4] utilize statistical features to model users’ overall language patterns and represent their long-term traits. Some approaches, like [11] concatenate users’ entire texts, while others, like [30] concatenate representations from different texts to learn the overall user representation. These methods primarily focus on capturing the consistent personality traits that persist over an extended period. However, they may not fully capture the dynamic changes and fluctuations in a user’s personality that can occur in short-term states.
On the other hand, the second category of methods adopts a hierarchical strategy, focusing on learning the current short-term representation of users’ short-term states using their current posts. These short-term representations are then aggregated to form an overall user representation for personality detection. For example, [18] employed a hierarchical attention network [34] and [20] employed hierarchical convolutional neural network (CNN) to generate personality representations from posts for personality detection. [33, 32] utilized graph neural networks to aggregate text representations. These approaches capture short-term variations in personality by modeling dependencies within each post and aggregating representations across multiple posts to create an overall user representation. However, while effective in capturing short-term variations and dependencies within each post, they may overlook the enduring traits that remain consistent over an extended period, thus not fully capturing the long-term stable aspects of personality.
@@
3 Methodology
3.1 Problem Definition
@@ 3.2 Overview of DEN
3.3 Input Layer
The Input Layer is responsible for acquiring the inputs for the Dual Personality Encoding Layer from user-generated posts $P=\{p_{1},p_{2},...,p_{N}\}$ and an external LIWC Psychological Knowledge Base [23]. Specificly, we need to prepare input for the Long-Term User Encoding module, which requires a psychological graph, denoted as $\mathcal{G}(M,A)$. This graph comprises a entity node feature matrix $M\in\mathbb{R}^{K\times d_{g}}$ representing entity embeddings, and an adjacency matrix $A\in\mathbb{R}^{K\times K}$ that captures the relationships between entity nodes. Here, $K$ represents the number of entities, and $d_{g}$ denotes the feature dimension.
To construct the psychological graph $\mathcal{G}(M,A)$, on the one hand, we need to obtain the feature of psychological entities for each user. These psychological entities, denoted as $E_{u}=\{w_{1},w_{2},...,w_{K}\}$, are defined as keywords mentioned by the user, and are automatically extracted from the user posts collections through the python tool provided by LIWC (Linguistic Inquiry and Word Count) [^1] https://github.com/chbrown/liwc-python. Then, we employ GloVe [24] to convert these entities into node features. We also experimented with using pretrained language model like BERT [13] for feature extraction, but there was no improvement in performance, and in fact, there was a noticeable decline in performance. We speculate that this is because there is no need to utilize the context to get the entity node features, which is the advantage of BERT. Formally, the feature vector for each entity node is acquired using GloVe word embeddings, as shown in Equation (1):
@@ 3.3 Input Layer
On the other hand, we need to establish edges between the acquired entity nodes based on the external LIWC Psychological Knowledge Base. Specifically, when two entities share the same LIWC category, we create an edge between them. This process results in an adjacency matrix $A\in\mathbb{R}^{K\times K}$, where $A_{ij}=1$ signifies an edge between node $i$ and $j$, and $A_{ij}=0$ indicates no connection. This process can be formulated as shown in Equation (2):
@@
3.4 Dual Personality Encoding Layer
Dual Personality Encoding Layer is the core of DEN, which enables us to learn the long-term stable user personality trait representations by analyzing the usage patterns of psychological entities based on the psychological graph. Simultaneously, it obtains the current short-term user personality state representations from the user’s various posts. This layer allows for a comprehensive understanding of the user’s personality by considering both the enduring traits and the transient states expressed in their posts.
@@ 3.4 Dual Personality Encoding Layer
3.4.1 Long-term Personality Encoding
@@
3.4.2 Short-term Personality Encoding
In the Short-term Personality Encoding module, the user’s short-term personality states is represented by each of the user’s post semantic representations. To obtain the contextual representation of each post, we utilize the pre-trained language model BERT [13] to acquire the contextual representation of each post $p_{i}$. The contextual representation of each post, denoted as $p_{i}$, is acquired as follows:
@@ 3.4.2 Short-term Personality Encoding
where $p_{i}$ represents the $i$-th post of the user, $\mathrm{BERT}(\cdot)$ refers to the pre-trained BERT’s hidden state of the “[CLS]” token in each post text, and $d$ denotes the hidden dimension of BERT. The “[CLS]” token in BERT is added as the first token of each post and is used to capture the overall content and context of the entire input text.
@@ 3.4.3 Bidirectional Interaction
Specifically, we first establish a bipartite graph, denoted as $\widetilde{\mathcal{G}}(\widetilde{M},\widetilde{A})$. The node matrix $\widetilde{M}\in\mathbb{R}^{(K+N)\times d}$ is the concatenation of long-term personality representations $M^{l}$ and short-term representations $H^{l}$. The adjacency matrix $\widetilde{A}$ represents the relationship between user posts and entities, with $\widetilde{A}_{ij}$ equaling 1 if entity $e_{i}$ appears in post $p_{j}$, and 0 otherwise.
Based on the bipartite $\widetilde{\mathcal{G}}(\widetilde{M},\widetilde{A})$, we utilize another GCN to refine the representation of the user’s long-term and short-term aspects, which is formalized as follows:
@@ 3.6 Objective Function
where $y_{t}\in\{0,1\}$ represents the ground truth personality characteristic label, and $\hat{y}_{t}\in\{0,1\}$ is the predicted logit for the $t$-th personality characteristic. The binary cross-entropy loss function is particularly well-suited for binary classification tasks like personality characteristic prediction. It enables the model to consider both the long-term stable personality of users and their dynamic changes over time, thereby improving the accuracy of personality detection based on user-generated content.
4 Experiments
4.1 Datasets
In the field of personality detection, the two most popular personality models are the Big Five [7] and MBTI (Myers-Briggs Type Indicator) [22]. Due to our psychological theoretical analysis being based on the MBTI personality model, and the ease of obtaining MBTI labels making it more applicable for developing a personality detection model toward social media [2], we have chosen to utilize two widely used MBTI personality detection datasets for our research. The MBTI personality model classifies personality types into four characteristics include Extroversion vs. Introversion (E/I), Sensing vs. Intuition (S/N), Thinking vs. Feeling (T/F), and Judging vs. Perception (J/P).
We utilized the publicly available Kaggle MBTI and Pandora MBTI datasets for our evaluations. The Kaggle dataset [^2] https://www.kaggle.com/datasets/datasnaek/mbti-type consists of 8675 users, each contributing approximately 45-50 posts. The Pandora dataset [^3] https://psy.takelab.fer.hr/datasets/all/ contains dozens to hundreds of posts for each of the 9067 users. Detailed statistics of the Kaggle and Pandora datasets are presented in 4.
From Figure 4, it is evident that both datasets exhibit varying degrees of distribution imbalance across different personality dimensions. Therefore, to evaluate the performance on each personality characteristic, we employed the Macro-F1 metric. And the average Macro-F1 of all the personality characteristics was used to assess the overall performance.
We shuffled the datasets and divided them into a 6:2:2 split for training, validation, and testing, respectively. Consistent with prior studies [11, 32], we removed all words matching any personality label to prevent information leaks.
4.2 Baseline Methods
@@
SVM [4] and XGBoost [26]: SVM or XGBoost is employed to classify personality characteristics based on the TF-IDF features extracted from the document, which is created by concatenating user-generated posts.
BiLSTM [27]: This model encodes each post using a Bi-LSTM model with GloVe word embeddings and obtains the user representation by averaging the post representations.
AttRCNN [29]: This model utilizes a CNN-based aggregator to obtain the user representation while incorporating psycholinguistic knowledge by concatenating Linguistic Inquiry and Word Count (LIWC) features.
BERT [12]: In this model, fine-tuned BERT is employed to encode each post, and the user representation is derived by mean pooling over the post representations.
SN + Attn [18]: This model utilizes a hierarchical attention network, employing a GRU with word-level attention to encode each post and another GRU with post-level attention to generate the user representation.
Trans-MD [30]: This model sequentially encodes the posts using BERT and stores them in memory, enabling posts to access the information of former ones.
TrigNet [33]: This model constructs a tripartite graph with three types of nodes: post, word, and category, and aggregates post information using a graph neural network.
D-DGCN [32]: D-DGCN is a dynamic graph neural network that learns correlations within user posts and aggregates them based on the learned graph structure.
Notably, D-DGCN employed a domain-adaptive pretrained BERT as its backbone. To ensure equitable experimental comparison, we adopt D-GCN as a baseline with the original pretrained BERT.
4.3 Implementation Details

### v2

@@ Abstract
Textual personality detection aims to identify personality characteristics by analyzing user-generated content on social media platforms. Extensive psychological literature highlights that personality encompasses both long-term stable traits and short-term dynamic states. However, existing studies often concentrate only on either long-term or short-term personality representations, neglecting the integration of both aspects. This limitation hinders a comprehensive understanding of individuals’ personalities, as both stable traits and dynamic states are vital. To bridge this gap, we propose a Dual Enhanced Network (DEN) to jointly model users’ long-term and short-term personality traits. In DEN, the Long-term Personality Encoding module models stable long-term personality traits by analyzing consistent patterns in the usage of psychological entities. The Short-term Personality Encoding module captures dynamic short-term personality states by modeling the contextual information of individual posts in real-time. The Bi-directional Interaction module integrates both aspects of personality, creating a cohesive and comprehensive representation of the user’s personality. Experimental results on two personality detection datasets demonstrate the effectiveness of the DEN model and underscore the importance of considering both stable and dynamic aspects of personality in textual personality detection.
I Introduction
Textual personality detection toward social media aims to identify an individual’s personality characteristics reflected in their written text [1].
By analyzing a user’s personality characteristics, we can gain a deeper understanding of their preferences [2], behavior patterns [3], and decision-making processes [4]. As a result, personality detection has been widely applied in various downstream tasks, such as personalized advertising and recommendation systems [5], dialogue systems [6], and individual career planning [7].
@@ I Introduction
Traditionally, personality assessment relies on questionnaires designed by psychologists, which can be time-consuming and labor-intensive, and impractical for the online world. The rise of social media has created vast opportunities for automated textual personality detection, attracting a growing number of researchers in the field [8].
Textual personality detection, although resembling a multi-label classification task, poses unique challenges. For example, the input comprises multiple posts rather than a single text [1]. Each post can reflect the user’s current short-term personality state, while the collection of all posts can capture the user’s long-term trait. To address these challenges, numerous studies have been conducted, broadly categorized into two main types.
On the one hand, some methods utilize a flat structure and focus on modeling the overall long-term representation of users to capture their stable personality traits for personality detection. This can be accomplished by utilizing various linguistic or statistical features [9, 10] or concatenating the user’s post representations into a single representation [11] to capture the long-term representation.
On the other hand, another line of research, employing a hierarchical structure, involves first modeling the current short-term representation of users’ states using their individual posts. Subsequently, these representations are aggregated to obtain an overall user representation for personality detection [12, 13, 1].
@@
Figure 2: Distribution of changes in MBTI personality characteristics among 84 college students over a 6-year period (from the first year of pre-university to the fourth year of undergraduate study). The bar chart compares the number of personality changes among different gender groups, while the line graph illustrates the trend of the percentage of different personality change counts. Data are from [18]
@@
The above findings emphasize the complexity of personality development, with a combination of stability and change occurring simultaneously. It underscores the importance of considering both the stable long-term traits and the dynamic short-term states when studying and predicting personality.
Unfortunately, it is non-trivial to capture both the long-term and short-term perspectives of user personality due to the following challenges: (1) Effectively modeling long-term personality representations, which reflect stable and consistent traits over time; (2) Accurately modeling short-term personality states, which are dynamic and context-dependent; and (3) Integrating these two distinct representations to enhance personality prediction.
To address these challenges, we propose the Dual Enhanced Network (DEN) for personality detection, which jointly captures long-term and short-term personality representations while dynamically integrating them to enhance personality detection. Specifically, DEN introduces three novel modules: (1) Long-term Personality Encoding: This module models users’ long-term stable personality traits by aggregating their consistent usage patterns of psychological cues over time. It achieves this by constructing a user-specific psychological graph using LIWC (Linguistic Inquiry and Word Count) [9]. By applying graph neural network (GNN) to this graph, this module effectively analyzes the relationships between psychological entities, learning insightful representations of enduring personality characteristics. (2) Short-term Personality Encoding: This module models users’ short-term dynamic personality states by capturing their immediate linguistic expressions in individual posts. Leveraging a pre-trained language model with multi-head self-attention mechanism, it encodes each post to reflect the user’s transient personality changes, enabling effective representation of short-term personality states. (3) Bi-directional Interaction: This module uses a bipartite heterogeneous graph to facilitate interaction between long-term and short-term representations. Through this interaction, both perspectives are mutually refined, resulting in a more comprehensive and accurate personality representation.
Extensive experiments demonstrate the effectiveness of DEN in the personality detection task. The key contributions of this paper can be summarized as follows:
@@ II Related Work
One type of research in textual personality detection utilizes a flat structure, focusing on modeling the overall long-term representation of users, which reflects their stable personality traits. Various approaches have been proposed in this category. For instance, studies such as [21, 22, 23] employ various linguistic features extracted from the entirety of users’ posts to capture their language habits and represent their long-term traits. Similarly, [10, 24] utilize statistical features to model users’ overall language patterns and represent their long-term traits. Some approaches, like [11] concatenate users’ entire texts, while others, like [13, 25] concatenate representations from different texts to learn the overall user representation. These methods primarily focus on capturing the consistent personality traits that persist over an extended period. However, they may not fully capture the dynamic changes and fluctuations in a user’s personality that can occur in short-term states.
On the other hand, the second category of methods adopts a hierarchical strategy, focusing on learning the current short-term representation of users’ short-term states using their current posts. These short-term representations are then aggregated to form an overall user representation for personality detection. For example, [12] employed a hierarchical attention network [26] and [27] employed hierarchical convolutional neural network (CNN) to generate personality representations from posts for personality detection. [28, 1, 29] utilized graph neural networks to aggregate text representations. These approaches capture short-term variations in personality by modeling dependencies within each post and aggregating representations across multiple posts to create an overall user representation. However, while effective in capturing short-term variations and dependencies within each post, they may overlook the enduring traits that remain consistent over an extended period, thus not fully capturing the long-term stable aspects of personality.
@@
III Methodology
III-A Problem Definition
@@ III-B Overview of DEN
III-C Input Layer
The Input Layer is responsible for acquiring the inputs for the Dual Personality Encoding Layer from user-generated posts $P=\{p_{1},p_{2},...,p_{N}\}$ and an external LIWC Psychological Knowledge Base [9]. Specifically, we need to prepare input for the Long-Term User Encoding module, which requires a psychological graph, denoted as $\mathcal{G}(M,A)$. This graph comprises a entity node feature matrix $M\in\mathbb{R}^{K\times d_{g}}$ representing entity embeddings, and an adjacency matrix $A\in\mathbb{R}^{K\times K}$ that captures the relationships between entity nodes. Here, $K$ represents the number of entities, and $d_{g}$ denotes the feature dimension. To construct the psychological graph $\mathcal{G}(M,A)$, we perform the following steps:
On the one hand, we extract the features of psychological entities for each user. These psychological entities, denoted as $E_{u}=\{w_{1},w_{2},...,w_{K}\}$, are defined as psychological keywords mentioned by the user, and are automatically extracted from the user posts collections through the python tool provided by LIWC (Linguistic Inquiry and Word Count, a tool that detects words indicative of psychological states including emotion, cognition, and social orientation) [^1] https://github.com/chbrown/liwc-python. These entities include words appearing in user posts that are categorized by LIWC. For example, in the sentence “I am happy today”, the word “happy” is identified as part of the LIWC category affect and is thus included as a psychological entity in the user’s psychological graph. Next, we employ GloVe [30] to convert these entities into node features.
Formally, the feature vector for each entity node is acquired using GloVe word embeddings, as shown in Equation (1):
@@ III-C Input Layer
On the other hand, we establish edges between the extracted entity nodes based on co-categories in the LIWC knowledge base to capture the relationships between them. Specifically, when two entities share at least one LIWC category, an edge is created between them. For example, “happy” and “sad” both belong to the category affect, so an edge is established between these nodes in the graph. This process results in an adjacency matrix $A\in\mathbb{R}^{K\times K}$, where $A_{ij}=1$ signifies an edge between node $i$ and $j$, and $A_{ij}=0$ indicates no connection. This process can be formulated as shown in Equation (2):
@@
III-D Dual Personality Encoding Layer
Dual Personality Encoding Layer is the core of DEN, which enables us to learn the long-term stable user personality trait representations by analyzing the usage patterns of psychological entities based on the psychological graph. Simultaneously, it obtains the current short-term user personality state representations from dynamic context of users’ individual posts. This layer allows for a comprehensive understanding of the user’s personality by considering both the enduring traits and the transient states expressed in their posts.
@@ III-D Dual Personality Encoding Layer
III-D1 Long-term Personality Encoding
@@
III-D2 Short-term Personality Encoding
In the Short-term Personality Encoding module, the user’s short-term personality states is represented by the dynamic context of users’ individual posts. To obtain the contextual representation of each post, we utilize the pre-trained language model BERT [32] to acquire the contextual representation of each post $p_{i}$. The contextual representation of each post, denoted as $p_{i}$, is acquired as follows:
@@ III-D2 Short-term Personality Encoding
where $p_{i}$ represents the $i$-th post of the user, $\mathrm{BERT}(\cdot)$ refers to the pre-trained BERT’s hidden state of the “[CLS]” token in each post text, and $d$ denotes the hidden dimension of BERT.
The “[CLS]” token in BERT is added as the first token of each post and is used to capture the overall content and context of the entire input text.
@@ III-D3 Bidirectional Interaction
[…]
@@ III-F Objective Function
where $y_{t}\in\{0,1\}$ represents the ground truth personality characteristic label, and $\hat{y}_{t}\in\{0,1\}$ is the predicted logits for the $t$-th personality characteristic. The binary cross-entropy loss function is particularly well-suited for binary classification tasks like personality characteristic prediction. It enables the model to consider both the long-term stable personality of users and their dynamic changes over time, thereby improving the accuracy of personality detection based on user-generated content.
IV Experiments
IV-A Datasets
In the field of personality detection, the two most popular personality models are the Big Five [34] and MBTI (Myers-Briggs Type Indicator) [19].
Due to our psychological theoretical analysis being based on the MBTI personality model, and the ease of obtaining MBTI labels making it more applicable for developing a personality detection model toward social media [35], we have chosen to utilize two widely used MBTI personality detection datasets for our research. The MBTI personality model classifies personality types into four characteristics include Extroversion vs. Introversion (E/I), Sensing vs. Intuition (S/N), Thinking vs. Feeling (T/F), and Judging vs. Perception (J/P).
We utilized the publicly available Kaggle MBTI and Pandora MBTI datasets for our evaluations. The Kaggle dataset [^2] https://www.kaggle.com/datasets/datasnaek/mbti-type consists of 8675 users, each contributing an average of 47 posts. The Pandora dataset [^3] https://psy.takelab.fer.hr/datasets/all/ contains 9067 users, each contributing an average of 96 posts.
We shuffled the datasets and divided them into a 6:2:2 split for training, validation, and testing, respectively. Consistent with prior studies [11, 1], we removed all words matching any personality label to prevent information leaks. Detailed statistics of the Kaggle and Pandora datasets are presented in 4.
From Figure 4, it is evident that both datasets exhibit varying degrees of distribution imbalance across different personality dimensions. Therefore, to evaluate the performance on each personality characteristic, we employed the Macro-F1 metric. And the average Macro-F1 of all the personality characteristics was used to assess the overall performance. Statistical significance of pairwise differences for DEN against the best baseline is determined by the t-test ($p<0.05$).
IV-B Baseline Methods
@@
SVM [24] and XGBoost [36]: Employs SVM or XGBoost to classify personality characteristics based on the TF-IDF features extracted from the document, which is created by concatenating user-generated posts.
BiLSTM [37]: Encodes each post using a Bi-LSTM model with GloVe word embeddings and obtains the user representation by averaging the post representations.
AttRCNN [20]: Utilizes a CNN-based aggregator to obtain the user representation while incorporating psycholinguistic knowledge by concatenating Linguistic Inquiry and Word Count (LIWC) features.
BERT [33]: Fine-tunes BERT to encode individual posts, with the user representation derived by mean pooling over the post representations.
SN + Attn [12]: Utilizes a hierarchical attention network, employing a GRU with word-level attention to encode each post and another GRU with post-level attention to generate the user representation.
Trans-MD [13]: Sequentially encodes the posts using BERT and stores them in memory, enabling posts to access the information of former ones.
TrigNet [28]: Constructs a tripartite graph with three types of nodes: post, word, and category, and aggregates post information using a graph neural network.
D-DGCN [1]: Employs a dynamic graph neural network to model correlations within user posts and aggregates these representations based on the graph structure.
LLMs: Use task-specific prompts [38] to guide large language models (including DeepSeek V3 and GPT-4o) in inferring user personality based on their posts.
PsyCoT [38]: Guides LLMs to simulate the process of completing psychological questionnaires through multi-turn dialogue for user personality prediction.
IV-C Implementation Details